Circular No. 6429

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1996 N1

Additional precise positions:

1996 UT α_{2000} δ_{2000} Observer
July 4.85856 11^{h}35^{m}28.^{s}28 +4^{o}42′27′′.8 Koleny
4.86485 11 35 29.38 +4 42 49.8  _{”}
4.87268 11 35 30.75 +4 43 17.1  _{”}
4.88944 11 35 33.54 +4 44 18.1 Ewen-Smith
4.89104 11 35 33.79 +4 44 23.0  _{”}
4.89288 11 35 34.12 +4 44 30.0  _{”}
4.89369 11 35 34.27 +4 44 32.6  _{”}
4.89455 11 35 34.58 +4 44 33.6  _{”}
4.89770 11 35 35.01 +4 44 45.4  _{”}
4.90010 11 35 35.40 +4 44 55.9  _{”}
4.90085 11 35 35.55 +4 44 56.9  _{”}

P. Koleny and L. Kornos (Modra). 0.6-m f/5.5 reflector + CCD. B. Ewen-Smith and D. Farley (Portimao). 0.5-m reflector + CCD.

SS LACERTAE P. B. Etzel, N. H. Volgenau, and Q. T. Nguyen, Department of As-

tronomy, San Diego State University, communicate: “We have observed, for the first time, a double-lined spectrum of SS Lac (BD +45^{o}3782) with the 1.0-m Cassegrain spectrograph at Mount Laguna Observatory. Schiller et al. (1991, BAAS 23, 879) discuss the cessation of eclipses in this system; their twelve spectra obtained during 1982–1984 revealed only single lines and no detectable radial-velocity variation. Schiller [1996, Astronomy 24(7), 46] recently proposed that the two components were disrupted by a third object and sent into trajectories perpendicular to our line of sight. Our spectroscopic observation provides a velocity difference of 174 km/s between the components on 1996 July 1.47 UT, corresponding to orbital phase 0.78 from the eclipse ephemeris of Dugan and Wright (1935, AJ 44, 150). Thus, SS Lac is probably still a binary system, which we are continuing to observe.”

R CORONAE BOREALIS Visual magnitude estimates by M. V. Zanotta, Milan, Italy (cf. IAUC

6260): Apr. 16.92, 12.2; May 21.09, 9.3; June 10.88, 8.5; July 3.85, 8.1.

1996 July 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green